### v1

# Symbolic Powers via Extension

Abstract.
This article investigates under which conditions the symbolic powers of the extension of an ideal is the same as the extension of the symbolic powers. Our result generalizes the known scenarios. As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is algebraically closed. Initially, these were known for ideals in polynomial rings.
Key words and phrases:
Resurgence, asymptotic resurgence, prime extension, flat, faithfully flat
2010 Mathematics Subject Classification:
14N20, 13F20, 14C20
1. Introduction
For an ideal $I$ in a noethrian ring $R$, the $m$-th symbolic power is defined as follows:
$I^{(m)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{m}R_{\mathfrak{p}}\cap R.$
The symbolic powers are very well studied objects. The behavior of symbolic powers under homomorphism are also well studied. If $\phi:A\to B$ is a homomorphism of rings, then for any ideal $I$ in $A$ we use $IB$ to denote the extension of the ideal $I$ in $B$ and for any ideal $J$ in $B$ we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$.
In [GHMN17, Proposition 3.8], it has been proved that if $\phi:A={\mathbbm{k}}[y_{1},\dots,y_{s}]\to B={\mathbbm{k}}[x_{0},\dots,x_{n}]$ is defined by $y_{i}\mapsto f_{i}$, with every subset of $c+1$ many $f_{i}$ forms $B$-regular sequence then $I^{(m)}B=\left(IB\right)^{(m)}$, where $I$ is the Stanley-Reisner ideal of a matroid on $[s]=\{1,\dots,s\}$ of dimension $s-1-c$. In [Ake17, Theorem 4], it has been proved that, if $\phi:A={\mathbbm{k}}[y_{0},\dots,y_{n}]\to B={\mathbbm{k}}[y_{0},\dots,y_{n}]$ is given by $y_{i}\to f_{i}$, where $f_{0},\dots,f_{n}$ is an $B$-regular sequence of homogeneous element of the same degree, then $(IB)^{(m)}=(I^{(m)})B$, where $I$ is the defining ideal of a 00-dimensional scheme. In [Wal18, Proposition 2.1], it has been proved that if $\phi:A\to B$ is faithfully flat and the associative primes of $I$ extends to associative primes of $IB$ then the symbolic powers of the extension is the extension of that of $I$.
Motivated by these, in this article, we investigate the following question.
Question 1.1.
Let $A$ and $B$ be noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(m)}=I^{(m)}B$, where $m\geqslant 2$?
In [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1], the key tool to prove symbolic power extends to symbolic power was that the morphisms were faithfully flat. We proved that one can relax faithfully flatness and prove something more general. We prove the following result extending the known situations where symbolic powers of the extension is the extension of the symbolic powers.
Main Result 1 (Theorem 2.5).
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(m)}=I^{(m)}B$.
In Remark 2.6, we discussed that our hypothesis is generalizes the known situations. Thus it covers more classes and will broader applications.
One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$. In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$.
In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals. The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
If $I$ is a homogeneous ideal in a finitely generated ${\mathbbm{k}}$-algebra domain $A$, then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
$\rho(I)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subseteq I^{r},m\geqslant 1,r\geqslant 1\right\},$
${\rho_{a}}(I)=\sup\left\{\frac{m}{r}~{}|~{}I^{(mt)}\not\subseteq{I^{rt}},m\geqslant 1,r\geqslant 1,\text{ and }t\gg 1\right\}.$
If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention.
Motivated by 1.1 we study the following question.
Question 1.2.
Let $A$ and $B$ be noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24]. As an application of Theorem 2.5 and Corollary 3.4, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
Main Result 2.
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes B$. Then the following hold .
  1. (1)
    (Theorem 4.5) $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
  2. (2)
    (Theorem 4.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$ Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
The paper is outlined as follows. In Section 2, we prove our main results regarding the relation symbolic powers of an ideal with its extension. In Section 3 we study prime extension property and its applications. In Section 4, we apply results from Section 2 and Section 3 we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis.
2. Symbolic Powers via flat extension
In this section we prove that with specific conditions, the symbolic powers of an extension is the extension of the symbolic powers. Before proceeding to the main theorem we need to prove the following lemma regarding flat morphisms.
Lemma 2.1.
Let $M$ be a $R$-module where $R$ is a commutative ring, and $I$ be an ideal of $R$, then $(R/I)\otimes_{R}M\cong M/IM$. Furthermore, if $M$ is flat over $A$, we also have $I\otimes_{R}M\cong IM$.
Proof.
The first part can be found in a standard book, e.g., Dummit and Foote. But we provide the proof for completion. After tensoring the exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$ by $M$, we get $I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. It’s a well-known fact that $R\otimes_{R}M\cong M$, then it induces a module homomorphism $\phi$ from $I\otimes M$ to $M$ defined by $\phi:I\otimes M\to M,\phi(r\otimes m)=rm$. Now the image of $\phi$ is $IM$. Since $R\otimes_{R}M\cong M$, the kernel of induced homomorphism from $M$ to $(R/I)\otimes M$ is exactly the image of $\phi$, then by isomorphism theorem, we have $M/IM\cong(R/I)\otimes M$. Since $M$ is a flat, $R$-module we get the following exact sequence: $0\rightarrow I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. Following the previous result, we immediately have $I\otimes_{R}M\cong IM$. ∎
Now we are ready to prove one of the key lemmas.
Lemma 2.2.
If a flat ring homomorphism $\phi:A\to B$ extends a prime $\mathfrak{p}$ to a prime $\mathfrak{p}B$, and $\mathfrak{p}$ is a contraction of some ideals in $B$, then $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$ as $A$-algebra.
Proof.
Applying Lemma 2.1 to $B_{\mathfrak{p}B}$, we have $B_{\mathfrak{p}B}\cong((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Our goal is to show $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$.
Define
$\psi_{1}:((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\rightarrow A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\text{ by }\psi_{1}\left(\frac{a\otimes b}{c\otimes d}\right)=\frac{a}{c}\otimes\frac{b}{d}.$
Similarly define
$\psi_{2}:A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\rightarrow((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\text{ by }\psi_{2}\left(\frac{a}{c}\otimes\frac{b}{d}\right)=\frac{a\otimes b}{c\otimes d}.$
First we need to show both maps are well-defined.
Since $c\otimes d\notin\mathfrak{p}\otimes_{A}B$, then $c\notin\mathfrak{p}$. We also can consider $c\otimes d=1\otimes cd\notin 1\otimes_{A}\mathfrak{p}B$, so $cd\notin\mathfrak{p}B$. We also know that $c\notin\mathfrak{p}$, this forces $d\notin\mathfrak{p}B$ since $\mathfrak{p}B$ is prime. Therefore $\psi_{1}$ is well-defined.
Now we prove that $\psi_{2}$ is also well-defined. Indeed, $c\notin\mathfrak{p}$ and $d\notin\mathfrak{p}B$, then $\phi(c)d\notin\mathfrak{p}B$ because $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. To see this, let $\mathfrak{p}=\mathfrak{q}\cap A$ for an ideal $\mathfrak{q}$ in $B$. From this we observe that $\mathfrak{p}B\cap A=\left(\left(\mathfrak{q}\cap A\right)B\right)\cap A=\mathfrak{q}\cap A={\mathfrak{p}}$. This implies that $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. Note that $\phi(A)\cap\mathfrak{p}B\supset\phi(\mathfrak{p})$. Let $y\in\phi(A)\cap{\mathfrak{p}}B$. Then $y=\phi(x)$ for some $x\in A$. We show that $x\in{\mathfrak{p}}$. Now since $y\in{\mathfrak{p}}B$ then $x\in{\mathfrak{p}}B\cap A={\mathfrak{p}}$.
Notice that $\psi_{1}$ and $\psi_{2}$ are inverses to each other.
Now we show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done. Verifying the map commutes with the multiplication is straightforward, we only verify the addition here. Let $\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}$ and $\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}$ be elements in $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Now
$\psi_{1}\left(\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}\right)+\psi_{1}\left(\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}\right)$ $=\frac{a_{1}}{c_{1}}\otimes\frac{b_{1}}{d_{1}}+\frac{a_{2}}{c_{2}}\otimes\frac{b_{2}}{d_{2}}$
$=\frac{a_{1}c_{2}}{c_{1}c_{2}}\otimes\frac{b_{1}d_{2}}{d_{1}d_{2}}+\frac{c_{1}a_{2}}{c_{1}c_{2}}\otimes\frac{d_{1}b_{2}}{d_{1}d_{2}}$
$=\left(\frac{1}{c_{1}c_{2}}\otimes\frac{1}{d_{1}d_{2}}\right)\cdot(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2}).$
Again,
$\psi_{1}\left(\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}+\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}\right)$ $=\psi_{1}\left(\frac{(a_{1}\otimes b_{1})\cdot(c_{2}\otimes d_{2})}{(c_{1}\otimes d_{1})\cdot(c_{2}\otimes d_{2})}+\frac{(a_{2}\otimes b_{2})\cdot(c_{1}\otimes d_{1})}{(c_{2}\otimes d_{2})\cdot(c_{1}\otimes d_{1})}\right)$
$=\psi_{1}\left(\frac{1}{c_{1}c_{2}\otimes d_{1}d_{2}}\cdot(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2})\right)$
$=\left(\frac{1}{c_{1}c_{2}}\otimes\frac{1}{d_{1}d_{2}}\right)\cdot\left(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2}\right).$
The following theorem determines the associated primes of an extension of an ideal.
Theorem 2.3.
[Mat89, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and Let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
$\operatorname{Ass}_{B}(E\otimes_{A}G)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(E)}\operatorname{Ass}_{B}(G/\mathfrak{p}G).$
Lemma 2.4.
Let $\phi:A\to B$ be a flat homomorphism. If $I\subset A$ is a ideal then we have $\operatorname{Ass}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}({\mathfrak{p}}B)$. Furthermore, if $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then we have $\operatorname{Ass}(I^{(n)}B)\subset\operatorname{Ass}(IB)$ for all $n\geqslant 1$.
Proof.
Since $\phi$ is a flat then $A/I\otimes_{A}B\cong B/IB$. Now by Theorem 2.3,
$\operatorname{Ass}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
Note that $\operatorname{Ass}(I^{(n)})\subset\operatorname{Ass}(I)$ for all $n\geqslant 1$. Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
$\operatorname{Ass}(I^{(n)}B)=\bigcup_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}\left(I^{(n)}\right)}\operatorname{Ass}({\mathfrak{p}}^{\prime}B)=\left\{{\mathfrak{p}}^{\prime}B~{}|~{}{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I^{(n)})\right\}\subset\operatorname{Ass}(IB),\text{ as }\operatorname{Ass}\left(I^{(n)}\right)\subset\operatorname{Ass}(I).$
The following proves that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
Theorem 2.5.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$ and every associated prime of $I$ is a contraction of some ideal in $B$ Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
Proof.
Using the definition of symbolic powers of an ideal and assumption, we get
$(IB)^{(n)}=\bigcap_{{{\mathfrak{p}}}\in\operatorname{Ass}_{B}(IB)}((IB)^{n}B_{{{\mathfrak{p}}}}\cap B)=\bigcap_{\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}}((IB)^{n}B_{{{\mathfrak{p}}}B}\cap B).$
Now apply Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
$(IB)^{n}B_{\mathfrak{p}B}\cap B$ $\cong(I^{n}\otimes_{A}B)(A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
$=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
(*) $\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}\cap B$
Now we prove the last isomorphism. Observe that $B=A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$. Applying a property from [Mat89, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$. Here we consider $M_{3}$ as $B_{\mathfrak{p}B}$, $M_{2}$ as $B$, and $M_{1}$ as $A$, then we can see $B_{\mathfrak{p}B}$ is flat over $A$. Now by [Mat89, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules. Then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}$, then we intersect it with $B$ as $A\otimes_{A}B\cong B$. Recall that $I^{(n)}=\bigcap_{\mathfrak{p}\in\operatorname{Ass}_{A}(I)}(I^{n}A_{\mathfrak{p}}\cap A)$. Therefore we have:
$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B$
[…]
(2.1) $=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{(n)}B_{{\mathfrak{p}}B}\cap B$
(2.2) $=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{(n)}BB_{{\mathfrak{p}}B}\cap B$
(2.3) $=I^{(n)}B_{W}\cap B,\text{ where }W=B-\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$
(2.4) $=(I^{(n)}B)B_{W}\cap B$
(2.5) $=I^{(n)}B$
First we prove equality 2.1. Since $\phi$ is flat, then using Lemma 2.1, we get
$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\left(\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\right)\otimes_{A}B_{{\mathfrak{p}}B}$
$=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\otimes_{A}B_{{\mathfrak{p}}B})=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
The equality 2.2 is straight forward.
The proof of equality 2.3 is similar to [Gri21, Remark 2.8]. We add the proof for completion. Note that $I^{(n)}B_{W}\cap B=\{r\in B~{}|~{}rs\in I^{(n)}\text{ for some }s\in W\}$ and thus $I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Now let $r\in\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Then for each ${\mathfrak{p}}B$ there is $s\not\in{\mathfrak{p}}B$ such that $rs\in I^{(n)}B$. Thus $I^{(n)}B:r$ is not contained in any ${\mathfrak{p}}B\in\operatorname{Ass}(IB)$. Since the associated primes are finite then by Prime avoidance theorem
$I^{(n)}B:r\not\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B.$
Thus there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [AM94, Proposition 3.11(ii)]
$(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s.$
Now by Lemma 2.4, we have $\operatorname{Ass}\left(I^{(n)}B\right)\subset\operatorname{Ass}(IB)$. Thus whenever $s\notin W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$. Thus by [AM94, Proposition 4.4(iii)] we get
$(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s=I^{(n)}B.$
Remark 2.6.
In Theorem 2.5 we extend [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1] by replacing the faithfully flat condition with weaker conditions. Note that faithfully flatness will implies the assumptions in Theorem 2.5. But not vice versa.
We give simple examples satisfying assumptions in Theorem 2.5 but not in [Wal18, proposition 2.1]. Let $A$ be a ring having at least two different prime ideals such that one is not contained in another one, and $B=A_{\mathfrak{p}}$ be localization of $A$ at an arbitrary prime $\mathfrak{p}$. It is a well-known fact that localization is an exact functor, so $B$ is a flat $A$-module, and prime ideal $\mathfrak{p}$ in $A$ is the contraction of $\mathfrak{p}B$. Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [AM94, ex 3.16].
One can take $A={\mathbb{Z}}$ and $B$ as $\mathbb{Z}_{(p)}$ where $p$ is a prime number. Take $I=(p^{m})$ for some $m\geqslant 1$. Then $I$ satisfies assumptions in Theorem 2.5. Furthermore, letting $B$ be any localization of $A$ will meet the assumptions in Theorem 2.5. This tells us that the stalk of $\operatorname{Spec}A$ gives us enough information of symbolic power of an ideal generated by the corresponding point in $\operatorname{Spec}A$.
Another example would be take $A={\mathbb{C}}[t]$ and $B={\mathbb{C}}[t]_{(p(t))}$, where $p(t)$ is an irreducible polynomial in ${\mathbb{C}}[t]$.
3. Resurgence and Prime extension
The following is from [Wal18, Proposition 2.1].
Proposition 3.1.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a faithfully flat morphism. If $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$, then $\rho(I)=\rho(IB)$.
Now we discuss some specific classes of morphisms which together with flatness gives faithfully flat. We start with the following definition.
Definition 3.2.
A ring homomorphism $\phi:A\to B$ has prime extension property if for every prime $\mathfrak{p}$ of $A$, $\phi({\mathfrak{p}})=\mathfrak{p}B$ is also prime in $B$. We also call $\phi$ is a prime extension.
Remark 3.3.
We recall some facts about the morphisms with prime extension properties.
  1. (1)
    In many literature, (for example, see, [HJ21]) prime extension has been defined as either ${\mathfrak{p}}B$ is a prime ideal or, the unit ideal. For us, prime extension means that prime ideals extends to a prime ideals.
  2. (2)
    Flat and prime extension implies faithfully flatness. Indeed, by [Mat89, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
Corollary 3.4.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. If $I\subset A$ is a proper ideal and $\phi$ has prime extension properties then $\rho(I)=\rho(IB)$ and $\rho_{a}(I)=\rho_{a}(IB)$.
Proof.
Since $\phi$ is prime extension then by Remark 3.3, it is faithfully flat. Thus using Theorem 2.5 we get that
$I^{(m)}\subset I^{r}\text{ if and only if }(IB)^{(m)}\subset(IB)^{r},\text{ for all }m,r.$
Hence
$\rho(I)=\rho(IB),\text{ and }\rho_{a}(I)=\rho_{a}(IB).$
Flat and prime extension has geometric meanings. Before we start to talk about them, we recall the definition.
Definition 3.5.
Let $f:X\to Y$ be a morphism of schemes, and let $y\in Y$ be a point. Let $\kappa(y)$ be the residue field of y, i.e. $\kappa(y)=\mathcal{O}_{y}/\mathfrak{m}_{y}$, and let $\operatorname{Spec}\,\kappa(y)\rightarrow Y$ be the natural morphism. Then we define the fiber of the morphism $f$ over the point $y$ to be the scheme
$X_{y}=X\times_{Y}\hskip 1.42262ptSpec\,\kappa(y).$
Note that fiber $X_{y}$ is a scheme over $\kappa(y)$.
Let $f:\operatorname{Spec}B\to\operatorname{Spec}A$ be a morphism of schemes, induced by ring homomorphism $\phi:A\to B$. Then the fiber over $\mathfrak{p}\in\operatorname{Spec}A$ is given by $\operatorname{Spec}B_{\mathfrak{p}}=\operatorname{Spec}\left(B\otimes_{A}\kappa(\mathfrak{p})\right)=\operatorname{Spec}\left(B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}\right)$. We have the following proposition:
Proposition 3.6.
If the ring homomorphism $\phi:A\to B$ is flat then $B_{\mathfrak{p}}$ is an integral domain if only if $\mathfrak{p}B$ prime.
Proof.
Since $\phi$ is flat, we can consider $\mathfrak{p}\otimes_{A}B$ as $\mathfrak{p}B$. It’s easy to see the followings:
$B_{\mathfrak{p}}\text{ is an integral domain }\iff B/\mathfrak{p}B\text{ is anintegral domain }\iff\mathfrak{p}B\text{ is prime}.$
This proposition implies that if $A$ is reduced, and $\phi$ has prime extension property, then $B$ is also reduced. Indeed, if $A$ contains the zero ideal as a prime ideal, so does $B$. It also gives us an enormous class of examples. If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain, which is geometrically interesting. Before we give an example, we recall the following lemma.
Lemma 3.7.
[Sta24, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
Example 3.8.
Let ${\mathbbm{k}}$ be an algebraically closed field, and let
$A={\mathbbm{k}}[t]\text{ and }B={\mathbbm{k}}[t,x,y]/(y-x^{2}-tx).$
Let $\phi$ to be the ring homomorphism from $A$ to $B$ by sending $t$ to $t$, and let $\Phi$ be the morphism from $\operatorname{Spec}B$ to $\operatorname{Spec}A$ induced by $\phi$. Note that $A\otimes_{A}B$ is a domain. It is not hard to see $\phi$ is a flat morphism. The map from $I\otimes_{A}B$ to $A\otimes_{A}B$ is always injective for every ideal $I\subset A$, then by Lemma 3.7 we get that $\phi$ is flat. The injectivity comes from the fact that both $y-x^{2}$ and $y-x^{2}-tx$ are irreducible in ${\mathbbm{k}}[t,x,y]$.
We identify closed points of $A$ with elements of ${\mathbbm{k}}$, and obviously the $(t-c)$ in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension. The fiber $\operatorname{Spec}B_{(t-c)}$ is given by the plane curve $y=x^{2}+cx$ in $\mathbb{A}_{\mathbbm{k}}^{2}$, which is an irreducible variety, and those plane curves are even homeomorphic to each other.
The situation doesn’t always perform so nicely. If we consider $A$ as ${\mathbbm{k}}[t,s]$ and $B$ as ${\mathbbm{k}}[t,s,x,y]/(y-x^{2}-tx)$, we also have $(t-c)$ as our point, then $B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ will be ${\mathbbm{k}}[s,s^{-1},x,y]/(y-x^{2}-cx)={\mathbbm{k}}(s)[x,y]/(y-x^{2}-cx)$. However we have $(t-a,s-b)$ as our point where $a$,$b\in{\mathbbm{k}}$, now the corresponding fiber is $\operatorname{Spec}{\mathbbm{k}}[x,y]/(y-x^{2}-ax)$, which is not homeomorphic to the previous one.
4. Resurgence of sum of ideals
Reviewing the recent research of resurgence number of sums of ideals for polynomial rings over fields, our goal is to prove these results for sums of finitely generated ${\mathbbm{k}}$-algebras which are integral domains where ${\mathbbm{k}}$ is algebraically closed. In [BHJT21, Theorem 2.6] it has been proved that if $I$ and $J$ are ideals in different polynomial rings then
$\rho_{a}(I+J)=\overline{\rho}_{\alpha}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}.$
In [KNT24, Theorem 2.3], it has been proved that
$\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
In this section we prove similar results for finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field. One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras. The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up. To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals.
Lemma 4.1.
Let $A$ and $B$ be finitely generated ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed, which are integral domains. If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes B$, then $\phi$ is flat and has prime extension property. Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$.
Proof.
Recall that the canonical ring homomorphism is defined by
$\phi:A\to A\otimes_{\mathbbm{k}}B\text{ defined by }a\rightarrow a\otimes_{\mathbbm{k}}1.$
Observe that if we can show $\phi$ is injective, then $I\rightarrow A\otimes_{A}R$ is injective for every ideal $I\subset A$. Hence by Lemma 3.7, $\phi$ is flat. Injectivity of $\phi$ is obvious since both $A$ and $B$ are fintely generated algebras over an algebraically closed field ${\mathbbm{k}}$, and integral domains. If there exists a nonzero element in ${\mathbbm{k}}$ annihilating some elements in $A$ (or in $B$), then this element has to be zero in $A$ (or in $B$), which is a contradiction.
Let $\mathfrak{p}\subset A$ be a prime ideal, we need to prove that ${\mathfrak{p}}R$ is also prime. Since $\phi$ is injective then $\mathfrak{p}R=\mathfrak{p}\otimes_{\mathbbm{k}}B$. Observe that $R/\mathfrak{p}R=\left(A/\mathfrak{p}\right)\otimes_{\mathbbm{k}}B$ because $R$ is an integral domain.
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Corollary 3.4. ∎
Remark 4.2.
To see why we need the conditions on $A$ and $B$, we will need some examples and theorems. Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain.
The next lemma also extends [BHJT21, Lemma 2.4]. The proof of [BHJT21, Lemma 2.4] depends on divisibility in the polynomial ring. While working with general domains, we treated them as vector spaces and used properties of basis.
Lemma 4.3.
Let $A$ and $B$ be finitely generated ${\mathbbm{k}}$-algebras, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ be the sum of ideal extensions of $I$ and $J$ in $R=A\otimes_{\mathbbm{k}}B$ then
$\rho_{a}^{\lim\sup}(I+J)\geqslant{\max}\left\{\rho_{a}^{{\lim\sup}}(I),\,\rho_{a}^{{\lim\sup}}(J)\right\}.$
Proof.
Take $\theta\in{\mathbb{Q}}$ such that $\theta>\rho_{a}^{\lim\sup}(I+J)$. Then there is a $s_{0}\in{\mathbb{N}}$ such that for all $t\geqslant s_{0},\theta>\rho(I+J,t)$, where $\rho(I+J,t)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subset I^{r},m\geqslant t,r\geqslant t\right\}$. Thus for any $h,r\in{\mathbb{N}}$ so that $h,r\geqslant t\geqslant s_{0}$ and $\dfrac{h}{r}\geqslant\theta$, one has $(I+J)^{r}\supset(I+J)^{(h)}\supset I^{(h)}$. Up to this point the proof idea is same as [BHJT21, Lemma 2.4]. Now we need to show that $I^{(h)}\subseteq I^{r}$.
Let $V$ be a ${\mathbbm{k}}$ basis of $I$, which extends to $V^{*}$ a ${\mathbbm{k}}$-basis of $A$ containing $1$. Let $W$ be a ${\mathbbm{k}}$-basis of $J$, which extends to $W^{*}$, a ${\mathbbm{k}}$-basis of $B$ containing $1$. Then $V\otimes W^{*}$ is a ${\mathbbm{k}}$-basis for $I\otimes B$ and $V^{*}\otimes W$ is a ${\mathbbm{k}}$-basis for $V\otimes W^{*}$. Also, $V^{*}\otimes W^{*}$ is a ${\mathbbm{k}}$-basis of $R=A\otimes B$. Let $f\in I^{(h)}$ then $f=\sum_{i=1}^{n}c_{i}v_{i}$, where $c_{i}\in{\mathbbm{k}}\setminus\{0\}$ and $v_{i}\in V$. Now
$f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(I+J)^{(h)}\subset(I+J)^{r}=\sum_{j=0}^{r}(IR)^{i}(JR)^{r-j}=\sum_{j=0}^{r}\left((IR)^{j}\cap(JR)^{r-j}\right).$
The last equality follows from [HNTT20, Lemma 3.1]. Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes_{k}1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$. Since $1\notin W$ then $v\otimes 1\notin A\otimes J$. Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $11\leqslant j\leqslant r-$ 1. Hence $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(IR)^{h}$. Now $(IR)^{h}\cong I^{h}\otimes B$ and thus $f\otimes 1\in(IR)^{h}$ implies that $f\in I^{h}$. This completes the proof. ∎
The following generalizes [BHJT21, Lemma 2.5] to finitely generated algebra domains. The proof of [BHJT21, Lemma 2.5] depends on the binomial summation formula and numerical inequalities, which holds in general domains as well, so we omit the proof.
Lemma 4.4.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ denotes $IR+JR$, where $R=A\otimes_{\mathbbm{k}}B$, then $\rho_{a}(I+J)\leqslant\max\left\{\rho_{a}^{\lim\sup}(I),\rho_{a}^{\lim\sup}(J)\right\}$.
Now we give the formula for the asymptotic resurgence of the sum of ideals in finitely generated ${\mathbbm{k}}$-algebra domians where ${\mathbbm{k}}$ is an algebraically closed field.
Theorem 4.5.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ denotes $IR+JR$, where $R=A\otimes_{\mathbbm{k}}B$, then
$\rho_{a}(I+J)=\overline{\rho}_{\alpha}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
Proof.
The proof has the same spirit as [BHJT21, Theorem 2.6] and it follows from Lemma 4.1, Lemma 4.3 and Lemma 4.4. ∎
Now we give the result regarding the summation formula for resurgence of ideals in finitely generated ideals in ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed field.
Theorem 4.6.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. Set $I+J=IR+JR$, where $R=A\otimes B$. Then
$\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
Proof.
We give a sketch of the proof as it follows the same pattern as [KNT24, Theorem 2.3]. First note that by Lemma 4.1, $\rho(IR)=\rho(I)$ and $\rho(JR)=\rho(J)$. The binomial summation formula for symbolic powers hold for ideals in noetherian ${\mathbbm{k}}$-algebras (see [HJKN23, Theorem 4.1]. The other numerical inequalities in [KNT24, Theorem 2.3] follows straightforward. Hence the required result follows. ∎

### v2

@@
 Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism.
-Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$.
-Then $(IB)^{(m)}=I^{(m)}B$.
+Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$.
+Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
 In Remark 2.6, we discussed that our hypothesis is generalizes the known situations. Thus it covers more classes and will broader applications.
@@
 The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
-If $I$ is a homogeneous ideal in a finitely generated ${\mathbbm{k}}$-algebra domain $A$, then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
+If $I$ is a homogeneous ideal in a finitely generated regular ${\mathbbm{k}}$-algebra domain $A$, where ${\mathbbm{k}}$ is perfect then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
 $\rho(I)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subseteq I^{r},m\geqslant 1,r\geqslant 1\right\},$
@@
 Question 1.2.
-Let $A$ and $B$ be noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
+Let $A$ and $B$ be regular noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
 The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24].
-As an application of Theorem 2.5 and Corollary 3.4, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
+As an application of Theorem 2.5 and Proposition 3.3, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
 Main Result 2.
-Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed.
+Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated regular ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed.
 Set $I+J=IR+JR$, where $R=A\otimes B$.
@@
 Acknowledgments.
-The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis.
+The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments.
+This is a part of the second author’s Ph.D.
+thesis.
+We are also thankful to Souvik Dey and Hop D.
+Nguyen for some helpful discussions.
 2. Symbolic Powers via flat extension
@@
 Let $\phi:A\to B$ be a flat homomorphism.
-If $I\subset A$ is a ideal then we have $\operatorname{Ass}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}({\mathfrak{p}}B)$.
-Furthermore, if $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then we have $\operatorname{Ass}(I^{(n)}B)\subset\operatorname{Ass}(IB)$ for all $n\geqslant 1$.
+We set $\operatorname{Ass}_{A}^{*}(I)=\cup_{n\in{\mathbb{N}}}\operatorname{Ass}_{A}(I^{n})$.
+If $I\subset A$ is a ideal then we have $\operatorname{Ass}_{B}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B)$.
+Furthermore, if $\operatorname{Ass}_{B}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}_{A}^{*}(I)\}$ then we have
+  1. (1)
+    $\operatorname{Ass}_{B}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)\}$,
+  2. (2)
+    $\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}_{B}(I^{(n)}B)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}_{A}(I)}{\mathfrak{p}}B\text{ for all }n\geqslant 1.$
 Proof.
 Since $\phi$ is a flat then $A/I\otimes_{A}B\cong B/IB$. Now by Theorem 2.3,
-$\operatorname{Ass}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
-Note that $\operatorname{Ass}(I^{(n)})\subset\operatorname{Ass}(I)$ for all $n\geqslant 1$.
+$\operatorname{Ass}_{B}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
+By the assumption ${\mathfrak{p}}B$ is prime for each ${\mathfrak{p}}\in\operatorname{Ass}_{A}(I)$. Hence we get
+$\operatorname{Ass}_{B}(IB)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}.$
+Note that by [HJKN23, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$.
+Thus we get $\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}\subset\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}$ for all $n\geqslant 1$.
 Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
-$\operatorname{Ass}(I^{(n)}B)=\bigcup_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}\left(I^{(n)}\right)}\operatorname{Ass}({\mathfrak{p}}^{\prime}B)=\left\{{\mathfrak{p}}^{\prime}B~{}|~{}{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I^{(n)})\right\}\subset\operatorname{Ass}(IB),\text{ as }\operatorname{Ass}\left(I^{(n)}\right)\subset\operatorname{Ass}(I).$
+$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
 The following proves that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
@@
 Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism.
-Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$ and every associated prime of $I$ is a contraction of some ideal in $B$ Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
+Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~{}|~{}\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$.
+Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
 Proof.
+First we check that as in Lemma 2.4, $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$.
 Using the definition of symbolic powers of an ideal and assumption, we get
@@
 $(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s.$
-Now by Lemma 2.4, we have $\operatorname{Ass}\left(I^{(n)}B\right)\subset\operatorname{Ass}(IB)$.
-Thus whenever $s\notin W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$.
+Now by Lemma 2.4, we have $\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$.
+Thus whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$.
 Thus by [AM94, Proposition 4.4(iii)] we get
@@
 3. Resurgence and Prime extension
-The following is from [Wal18, Proposition 2.1].
-Proposition 3.1.
-Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a faithfully flat morphism.
-If $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$, then $\rho(I)=\rho(IB)$.
-Now we discuss some specific classes of morphisms which together with flatness gives faithfully flat. We start with the following definition.
 Definition 3.1.
@@
 Indeed, by [Mat89, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
-Corollary 3.4.
+Proposition 3.3.
 Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism.
@@
 Proof.
-Since $\phi$ is prime extension then by Remark 3.3, it is faithfully flat. Thus using Theorem 2.5 we get that
+Since $\phi$ is prime extension then by remark 3.2, it is faithfully flat. Thus using Theorem 2.5 and [AM94, ex 3.16(I)] we get that
 $I^{(m)}\subset I^{r}\text{ if and only if }(IB)^{(m)}\subset(IB)^{r},\text{ for all }m,r.$
@@
 $\rho(I)=\rho(IB),\text{ and }\rho_{a}(I)=\rho_{a}(IB).$
+We want to mention that [Wal18, Proposition 2.1] also proved the similar result.
 Flat and prime extension has geometric meanings. Before we start to talk about them, we recall the definition.
@@
 $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
-In this section we prove similar results for finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field.
+In this section we prove similar results for finitely generated regular ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field.
 One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras.
@@
 In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals.
+In this section, we assume that all things are regular.
 Lemma 4.1.
@@
 It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property.
-The last equality follows from Corollary 3.4.
+The last equality follows from Proposition 3.3.
 ∎
@@
 It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain.
+Also note by [TY03, Theorem 6(e)] $A\otimes B$ is regular.
 The next lemma also extends [BHJT21, Lemma 2.4].

### v3

@@
 Abstract.
-This article investigates under which conditions the symbolic powers of the extension of an ideal is the same as the extension of the symbolic powers.
-Our result generalizes the known scenarios.
-As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is algebraically closed.
-Initially, these were known for ideals in polynomial rings.
+This article investigates under which conditions symbolic powers of the extension of an ideal are the same as the extension of the symbolic powers.
+As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated graded $\mathbbm{k}$}-algebras which are domains, where $\mathbbm{k}$} is algebraically closed.
+Initially, these were known for homogeneous ideals in polynomial rings.
 Key words and phrases:
@@
 $I^{(m)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{m}R_{\mathfrak{p}}\cap R.$
-The symbolic powers are very well studied objects.
-The behavior of symbolic powers under homomorphism are also well studied.
-If $\phi:A\to B$ is a homomorphism of rings, then for any ideal $I$ in $A$ we use $IB$ to denote the extension of the ideal $I$ in $B$ and for any ideal $J$ in $B$ we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$.
-In [GHMN17, Proposition 3.8], it has been proved that if $\phi:A={\mathbbm{k}}[y_{1},\dots,y_{s}]\to B={\mathbbm{k}}[x_{0},\dots,x_{n}]$ is defined by $y_{i}\mapsto f_{i}$, with every subset of $c+1$ many $f_{i}$ forms $B$-regular sequence then $I^{(m)}B=\left(IB\right)^{(m)}$, where $I$ is the Stanley-Reisner ideal of a matroid on $[s]=\{1,\dots,s\}$ of dimension $s-1-c$.
-In [Ake17, Theorem 4], it has been proved that, if $\phi:A={\mathbbm{k}}[y_{0},\dots,y_{n}]\to B={\mathbbm{k}}[y_{0},\dots,y_{n}]$ is given by $y_{i}\to f_{i}$, where $f_{0},\dots,f_{n}$ is an $B$-regular sequence of homogeneous element of the same degree, then $(IB)^{(m)}=(I^{(m)})B$, where $I$ is the defining ideal of a 00-dimensional scheme.
-In [Wal18, Proposition 2.1], it has been proved that if $\phi:A\to B$ is faithfully flat and the associative primes of $I$ extends to associative primes of $IB$ then the symbolic powers of the extension is the extension of that of $I$.
-Motivated by these, in this article, we investigate the following question.
+The symbolic powers are very well-studied objects for commutative algebra and algebraic geometry, especially for the multiplicity theory.
+We are interested in the behavior of symbolic powers under ring homomorphisms.
+If $\phi:A\to B$ is a homomorphism of rings, then for an ideal $I$ in $A$, we use $IB$ to denote the extension of the ideal $I$ in $B$, and for an ideal $J$ in $B$, we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$.
+In this article, we primarily focus on the following question.
 Question 1.1.
-Let $A$ and $B$ be noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(m)}=I^{(m)}B$, where $m\geqslant 2$?
-In [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1], the key tool to prove symbolic power extends to symbolic power was that the morphisms were faithfully flat.
-We proved that one can relax faithfully flatness and prove something more general.
-We prove the following result extending the known situations where symbolic powers of the extension is the extension of the symbolic powers.
+Let $A$ and $B$ be Noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(n)}=I^{(n)}B$, where $m\geqslant 1$?
+If the map $\phi$ is faithfully flat, then the Question 1.1 has been well studied, for examples check, [GHMNmatroid, Theorem 3.6], [Akessesh, Theorem 4], and [walker2018uniform, Proposition 2.1].
+Our result generalizes from faithfully flat map to flat map, which is the following:
 Main Result 1 (Theorem 2.5).
@@
 Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
-In Remark 2.6, we discussed that our hypothesis is generalizes the known situations. Thus it covers more classes and will broader applications.
-One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$.
-In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$.
-In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals.
-The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
-If $I$ is a homogeneous ideal in a finitely generated regular ${\mathbbm{k}}$-algebra domain $A$, where ${\mathbbm{k}}$ is perfect then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
+We also give examples that satisfy assumptions in Theorem 2.5 but not under the faithfully flat map, such as Example 2.8 and Example 2.9.
+Moreover, In Example 2.7, we show that the condition that associated primes are contractions is necessary.
+One of the central topics of symbolic powers is the containment between the symbolic and ordinary powers of ideals.
+The containment problem asks to find $m$ and $r$ for which $I^{(m)}\subset I^{r}$.
+In [EinLazarsfeldSmith, HochsterHuneke, MaScwede], it has been proved that if $I$ is a proper ideal of big height $h$ (the largest height of associated primes of $I$) in a regular ring containing a field, then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$.
+After this celebrated result, one can keep investigating exact numbers for containments.
+The resurgence, defined by Bocci and Harbourne [bocci2010comparing], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [guardo2013asymptotic] measure containment in more refined settings.
+If $I$ is a proper ideal in a Notherian ring $R$, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
 $\rho(I)=\sup\left\{\dfrac{m}{r}~|~I^{(m)}\not\subseteq I^{r},m\geqslant 1,r\geqslant 1\right\},$
 ${\rho_{a}}(I)=\sup\left\{\frac{m}{r}~|~I^{(mt)}\not\subseteq{I^{rt}},m\geqslant 1,r\geqslant 1,\text{ and }t\gg 1\right\}.$
-If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention.
-Motivated by 1.1 we study the following question.
+Motivated by Question 1.1 and resurgence numbers, we study the following question.
 Question 1.2.
-Let $A$ and $B$ be regular noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
-The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24].
-As an application of Theorem 2.5 and Proposition 3.3, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
+Let $A$ and $B$ be Notherian and regular $\mathbbm{k}$}-algebras, II be an ideal of $A$, and $J$ be an ideal of $B$.
+What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$ and $R=A\otimes_{{\mathbbm{k}}}B$?
+The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [fiberprojective, sharpbound].
+As an application of Theorem 2.5, we extend those results ideals in standard graded finitely generated $\mathbbm{k}$}-algebras which are domains, and $\mathbbm{k}$} is algebraically closed.
 Main Result 2.
-Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated regular ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed.
-Set $I+J=IR+JR$, where $R=A\otimes B$.
+Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated graded regular $\mathbbm{k}$}-algebra domains $A$ and $B$, where $\mathbbm{k}$} is algebraically closed.
+Set $I+J=IR+JR$, where $R=A\otimes_{{\mathbbm{k}}}B$.
 Then the following hold.
   1. (1)
-    (Theorem 4.5) $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
+    (Theorem 3.5) $\rho_{a}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}$, and also
+    $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)$, where $\bar{\rho}$ and $\bar{\rho_{\alpha}}$ are resurgence (asymptotic resurgence respectively) of integral closure.
   2. (2)
-    (Theorem 4.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$ Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
+    (Theorem 3.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
+    If $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
 The paper is outlined as follows.
-In Section 2, we prove our main results regarding the relation symbolic powers of an ideal with its extension.
-In Section 3 we study prime extension property and its applications.
-In Section 4, we apply results from Section 2 and Section 3 we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
+In Section 2, we prove our main result regarding the relation symbolic powers of an ideal with its extension, and we also study prime extension property and its applications.
+In Section 3, we apply results from Section 2 and we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
 Acknowledgments.
-The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments.
-This is a part of the second author’s Ph.D.
-thesis.
-We are also thankful to Souvik Dey and Hop D.
-Nguyen for some helpful discussions.
-2. Symbolic Powers via flat extension
-In this section we prove that with specific conditions, the symbolic powers of an extension is the extension of the symbolic powers.
-Before proceeding to the main theorem we need to prove the following lemma regarding flat morphisms.
+The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. We are also thankful to Souvik Dey and Hop D. Nguyen for some helpful discussions.
+2. Symbolic Powers and Resurgence via flat extension
+In this section we will show our main theorem (Theorem 2.5). Before proceeding to the main theorem we provide some useful lemmas. The first lemma is very well-known and we state it without proof.
 Lemma 2.1.
 Let $M$ be a $R$-module where $R$ is a commutative ring, and $I$ be an ideal of $R$, then $(R/I)\otimes_{R}M\cong M/IM$. Furthermore, if $M$ is flat over $A$, we also have $I\otimes_{R}M\cong IM$.
-Proof.
-The first part can be found in a standard book, e.g., Dummit and Foote.
-But we provide the proof for completion.
-After tensoring the exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$ by $M$, we get $I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$.
-It’s a well-known fact that $R\otimes_{R}M\cong M$, then it induces a module homomorphism $\phi$ from $I\otimes M$ to $M$ defined by $\phi:I\otimes M\to M,\phi(r\otimes m)=rm$.
-Now the image of $\phi$ is $IM$.
-Since $R\otimes_{R}M\cong M$, the kernel of induced homomorphism from $M$ to $(R/I)\otimes M$ is exactly the image of $\phi$, then by isomorphism theorem, we have $M/IM\cong(R/I)\otimes M$.
-Since $M$ is a flat, $R$-module we get the following exact sequence: $0\rightarrow I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$.
-Following the previous result, we immediately have $I\otimes_{R}M\cong IM$.
-∎
-Now we are ready to prove one of the key lemmas.
+The next result is also a standard one. Since we could not find a resource, we provide a complete proof.
 Lemma 2.2.
@@
 Therefore $\psi_{1}$ is well-defined.
-Now we prove that $\psi_{2}$ is also well-defined.
+We prove that $\psi_{2}$ is also well-defined.
 Indeed, $c\notin\mathfrak{p}$ and $d\notin\mathfrak{p}B$, then $\phi(c)d\notin\mathfrak{p}B$ because $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$.
@@
 Notice that $\psi_{1}$ and $\psi_{2}$ are inverses to each other.
-Now we show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done.
+Thereofre we just need to show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done.
 Verifying the map commutes with the multiplication is straightforward, we only verify the addition here.
 Let $\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}$ and $\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}$ be elements in $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$.
-Now
+Then
 $\psi_{1}\left(\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}\right)+\psi_{1}\left(\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}\right)$ $=\frac{a_{1}}{c_{1}}\otimes\frac{b_{1}}{d_{1}}+\frac{a_{2}}{c_{2}}\otimes\frac{b_{2}}{d_{2}}$
@@
 Theorem 2.3.
-[Mat89, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and Let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
+[matsumura1989commutative, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
 $\operatorname{Ass}_{B}(E\otimes_{A}G)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(E)}\operatorname{Ass}_{B}(G/\mathfrak{p}G).$
@@
 $\operatorname{Ass}_{B}(IB)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B)=\{{\mathfrak{p}}B~|~{\mathfrak{p}}\in\operatorname{Ass}(I)\}.$
-Note that by [HJKN23, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$.
+Note that by [ha2023binomial, Lemma 2.2] if ${\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})$ then ${\mathfrak{p}}\in\operatorname{Ass}(I^{n})$ and ${\mathfrak{p}}\subset{\mathfrak{q}}$ for some ${\mathfrak{q}}\in\operatorname{Ass}(I)$.
 Thus we get $\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}\subset\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}$ for all $n\geqslant 1$.
-Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
-$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
-The following proves that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
+Since $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~|~{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$ then using Theorem 2.3 we get
+$\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(I^{(n)}B)}{\mathfrak{q}}=\bigcup_{\begin{subarray}{c}{\mathfrak{q}}\in\operatorname{Ass}({\mathfrak{p}}B),\\ {\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)\end{subarray}}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I^{(n)})}{\mathfrak{p}}B\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B,\text{ as }\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}\left(I^{(n)}\right)}{\mathfrak{p}}\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}.$
+We are ready to prove the main theorem of this section.
+The theorem states that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal, then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
 Theorem 2.5.
@@
 Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
+Furthermore, if $I^{l}$ is a contraction of some ideals for all $\ell\geqslant 1$, then $I^{(m)}\subset I^{r}$, if and only if $IB^{(m)}\subset IB^{r}$, for all $m,r\geqslant 1$, then we get $\rho(IB)=\rho(I)$ and $\rho_{a}(IB)=\rho(I)$.
 Proof.
 First we check that as in Lemma 2.4, $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~|~{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$.
-Using the definition of symbolic powers of an ideal and assumption, we get
+Using the definition of symbolic powers of an ideal and the assumption, we get
 $(IB)^{(n)}=\bigcap_{{{\mathfrak{p}}}\in\operatorname{Ass}_{B}(IB)}((IB)^{n}B_{{{\mathfrak{p}}}}\cap B)=\bigcap_{\{{{\mathfrak{p}}}B~|~{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}}((IB)^{n}B_{{{\mathfrak{p}}}B}\cap B).$
-Now apply Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
+Applying Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
 $(IB)^{n}B_{\mathfrak{p}B}\cap B$ $\cong(I^{n}\otimes_{A}B)(A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
-$=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
-(*) $\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}\cap B$
-Now we prove the last isomorphism.
-Observe that $B=A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$.
-Applying a property from [Mat89, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$.
+Next we want to show that $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$.
+It is known that $A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$ if $S=B\backslash\mathfrak{p}B$ has no zero divisors.
+Suppose $S$ has no zero divisors, then applying a property from [matsumura1989commutative, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$.
 Here we consider $M_{3}$ as $B_{\mathfrak{p}B}$, $M_{2}$ as $B$, and $M_{1}$ as $A$, then we can see $B_{\mathfrak{p}B}$ is flat over $A$.
-Now by [Mat89, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules.
-Then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}$, then we intersect it with $B$ as $A\otimes_{A}B\cong B$.
-Recall that $I^{(n)}=\bigcap_{\mathfrak{p}\in\operatorname{Ass}_{A}(I)}(I^{n}A_{\mathfrak{p}}\cap A)$.
-Therefore we have:
-$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B$
-[…]
+By [matsumura1989commutative, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules, then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}$.
+Therefore, intersecting it with $A\otimes_{A}B$ will not change the original intersection.
+If $S$ has zero divisors, then $B\backslash B_{\mathfrak{p}B}$ has exactly the same set of zero divisors in $S$, denoted as $S^{\prime}$.
+Note that any $a\otimes b\in I^{n}A_{\mathfrak{p}}\otimes_{A}S^{\prime}$ is always 0 in $I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$, and similarly for $a\otimes b\in A\otimes_{A}S^{\prime}$ in $A\otimes_{A}B_{\mathfrak{p}B}$.
+Observe that these 0’s in $A\otimes_{A}B_{\mathfrak{p}B}$ are 0’s in $(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$, so we can always intersect $I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$ with $A\otimes_{A}B_{\mathfrak{p}B}$ first, then $A\otimes_{A}B$, which leads to the same equality.
+Therefore $(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)=((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$ as we desire.
+Applying lemma 2.1 again on $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$, we will have our second isomorphism $((I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)\cong(I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B}\cap B$.
+Note that lemma 2.1 is used previously for the “backward” direction for the first isomorphism, and now applying it again for the “forward” direction for our second isomorphism, so we will have a new equality (two isomorphisms give an identity map), which is $(IB)^{n}B_{\mathfrak{p}B}\cap B=(I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B}\cap B$.
+Then the following equality is given:
+$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}((I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B})\cap B$
+We want to prove the following labeled equalities:
+$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}((I^{n}A_{p}\cap A)B_{{\mathfrak{p}}B})\cap B$
 (2.1) $=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{(n)}B_{{\mathfrak{p}}B}\cap B$
@@
 (2.5) $=I^{(n)}B$
-First we prove equality 2.1. Since $\phi$ is flat, then using Lemma 2.1, we get
-$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\left(\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\right)\otimes_{A}B_{{\mathfrak{p}}B}$
-$=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\otimes_{A}B_{{\mathfrak{p}}B})=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
+First we prove equality 2.1. Since $\phi$ is flat, and extension commutes with intersection under flat morphism, we get
+$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
 The equality 2.2 is straight forward.
-The proof of equality 2.3 is similar to [Gri21, Remark 2.8].
-We add the proof for completion.
-Note that $I^{(n)}B_{W}\cap B=\{r\in B~{}|~{}rs\in I^{(n)}\text{ for some }s\in W\}$ and thus $I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$.
+The proof of equality 2.3 is similar to [grifo2021symbolic, Remark 2.8]. We add the proof for completion. Note that $I^{(n)}B_{W}\cap B=\{r\in B~|~rs\in I^{(n)}\text{ for some }s\in W\}$ and thus
+$\displaystyle I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right).$
 Now let $r\in\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$.
@@
 $I^{(n)}B:r\not\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B.$
-Thus there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
-Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [AM94, Proposition 3.11(ii)]
+Then there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
+Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [atiyah1994introduction, Proposition 3.11(ii)]
 $(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s.$
-Now by Lemma 2.4, we have $\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$.
-Thus whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$.
-Thus by [AM94, Proposition 4.4(iii)] we get
+Now by Lemma 2.4, we have
+$\displaystyle\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B.$
+Therefore whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where $\mathfrak{q}$} is an associated prime of $I^{(n)}B$. By [atiyah1994introduction, Proposition 4.4(iii)] we get
 $(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s=I^{(n)}B.$
+Note that $I^{(m)}\subset I^{(m)}B\cap A\subset IB^{r}\cap A=I^{r}B\cap A=I^{r}$, the last equality holds since $I^{r}$ is a contraction of some ideal. ∎
 Remark 2.6.
-In Theorem 2.5 we extend [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1] by replacing the faithfully flat condition with weaker conditions.
-Note that faithfully flatness will implies the assumptions in Theorem 2.5.
-But not vice versa.
-We give simple examples satisfying assumptions in Theorem 2.5 but not in [Wal18, proposition 2.1].
+Note that if placing $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~|~\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$ by $\operatorname{Ass}(IB^{N})=\{{\mathfrak{p}}B~|~\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}(I^{N})\}$, then the theorem 2.5 is only true up for $1\leqslant n\leqslant N$.
+If one uses the minimal prime definition of the symbolic powers, then it is enough to assume that $\operatorname{Min}(IB)=\{{\mathfrak{p}}B~|~{\mathfrak{p}}\in\operatorname{Min}(I)\}$.
+The following example shows that the condition in Theorem 2.5 that the associated primes are contraction of the ideals is necessary.
+Example 2.7.
+Consider the natural map $\phi:{\mathbb{Z}}[x]\to{\mathbb{Q}}[x]$.
+Note that it is a flat map.
+Take $I=\langle x,2\rangle\subset{\mathbb{Z}}[x]$.
+Then $I$ is not a contraction of any ideals of ${\mathbb{Q}}[x]$.
+Since $I$ is a maximal ideal, then $I^{(n)}=I^{n}$ for all $n\geqslant 1$.
+But $I{\mathbb{Q}}[x]=\langle x\rangle$.
+Hence $(I{\mathbb{Q}}[x])^{(n)}=\langle x^{n}\rangle$ for all $n\geqslant 1$.
+But $I^{(n)}{\mathbb{Q}}[x]=\langle x\rangle$ for all $n\geqslant 1$.
+Thus $(I{\mathbb{Q}}[x])^{(n)}\neq I^{(n)}{\mathbb{Q}}[x]$ for all $n\geqslant 2$.
+In Theorem 2.5 we extend [GHMNmatroid, Theorem 3.6], [Akessesh, Theorem 4], and [walker2018uniform, proposition 2.1] by replacing the faithfully flat condition with flat condition.
+Note that assumptions in Theorem 2.5 implies all these theorems and proposition, but not vice versa.
+We now give simple examples satisfying assumptions in Theorem 2.5, but violating the faithfully flat condition required for all theorems and proposition we just mentioned.
+Example 2.8.
+Consider the map $\phi:{\mathbb{Z}}\to{\mathbb{Z}}\left[1/\penalty 50p\right]$, where $p$ is a prime number, then this map is obviously flat, since $I\otimes_{{\mathbb{Z}}}{\mathbb{Z}}\left[1/\penalty 50p\right]\to{\mathbb{Z}}\otimes_{{\mathbb{Z}}}{\mathbb{Z}}\left[1/\penalty 50p\right]$ is always injective for any ideal $I\subset{\mathbb{Z}}$.
+However, it is not faithfully flat since $Spec({\mathbb{Z}}\left[1/\penalty 50p\right])\to Spec({\mathbb{Z}})$ is not surjective for closed points, in particular, $(p)$ is the closed point.
+Finding any ideal $I$ in $\mathbb{Z}$} such that $(p)$ is not an associated prime of $I$, then $I^{(n)}{\mathbb{Z}}\left[1/\penalty 50p\right]=(I{\mathbb{Z}}\left[1/\penalty 50p\right])^{(n)}$.
+This example is also a good comparison to Example 2.7, since any primes in $\mathbb{Z}$} is a contraction of the same primes in ${\mathbb{Z}}\left[1/\penalty 50p\right]$ except $(p)$.
+This example can be extended to PID and its prime ideal, since prime ideal (other than zero ideal) is also a maximal ideal, then it can be shown that map from PID to itself with reversing a prime ideal is flat but not faithfully flat by the same reason.
+Example 2.9.
 Let $A$ be a ring having at least two different prime ideals such that one is not contained in another one, and $B=A_{\mathfrak{p}}$ be localization of $A$ at an arbitrary prime $\mathfrak{p}$.
 It is a well-known fact that localization is an exact functor, so $B$ is a flat $A$-module, and prime ideal $\mathfrak{p}$ in $A$ is the contraction of $\mathfrak{p}B$.
-Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [AM94, ex 3.16].
+Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [atiyah1994introduction, ex 3.16].
 One can take $A={\mathbb{Z}}$ and $B$ as $\mathbb{Z}_{(p)}$ where $p$ is a prime number.
@@
 Another example would be take $A={\mathbb{C}}[t]$ and $B={\mathbb{C}}[t]_{(p(t))}$, where $p(t)$ is an irreducible polynomial in ${\mathbb{C}}[t]$.
-3. Resurgence and Prime extension
-Definition 3.1.
-A ring homomorphism $\phi:A\to B$ has prime extension property if for every prime $\mathfrak{p}$ of $A$, $\phi({\mathfrak{p}})=\mathfrak{p}B$ is also prime in $B$.
-We also call $\phi$ is a prime extension.
-Remark 3.2.
-We recall some facts about the morphisms with prime extension properties.
-  1. (1)
-    In many literature, (for example, see, [HJ21]) prime extension has been defined as either ${\mathfrak{p}}B$ is a prime ideal or, the unit ideal.
-For us, prime extension means that prime ideals extends to a prime ideals.
-  2. (2)
-    Flat and prime extension implies faithfully flatness.
-Indeed, by [Mat89, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
-Proposition 3.3.
-Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism.
-If $I\subset A$ is a proper ideal and $\phi$ has prime extension properties then $\rho(I)=\rho(IB)$ and $\rho_{a}(I)=\rho_{a}(IB)$.
+Applying our Theorem 2.5, we will have the following corollary, which is crucial for the main theorems in our section 3.
+We also explain the importance of this corollary in details at the beginning of the section 3.
+Corollary 2.10.
+Let $\phi:A\to B$ be a flat morphism. If ${\mathfrak{p}}B$ is a prime ideal for all primes $\mathfrak{p}$} in $A$, then $\rho(I)=\rho(IB)$, and $\rho_{a}(I)=\rho_{a}(I)$ for any proper ideal $I$.
 Proof.
-Since $\phi$ is prime extension then by remark 3.2, it is faithfully flat. Thus using Theorem 2.5 and [AM94, ex 3.16(I)] we get that
-$I^{(m)}\subset I^{r}\text{ if and only if }(IB)^{(m)}\subset(IB)^{r},\text{ for all }m,r.$
-Hence
-$\rho(I)=\rho(IB),\text{ and }\rho_{a}(I)=\rho_{a}(IB).$
-We want to mention that [Wal18, Proposition 2.1] also proved the similar result.
-Flat and prime extension has geometric meanings. Before we start to talk about them, we recall the definition.
+The proof follows directly from Theorem 2.5.
+Note that if ${\mathfrak{p}}B$ is prime for all primes $\mathfrak{p}$} in $A$, then $\phi$ is faithfully flat and $IB\cap A=I$ for all ideals $I$ in $A$.
+Indeed, by [matsumura1989commutative, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
+∎
+Morphisms above with the property that that extend every prime ideal as an prime ideal, also known as morphims with prime extension property, see [hochsterprime].
+We will explain the geometric meaning of flat and prime extension, but before that, we need to recall the definition of fiber over points first.
 Definition 2.11.
@@
 Since $\phi$ is flat, we can consider $\mathfrak{p}\otimes_{A}B$ as $\mathfrak{p}B$. It’s easy to see the followings:
-$B_{\mathfrak{p}}\text{ is an integral domain }\iff B/\mathfrak{p}B\text{ is anintegral domain }\iff\mathfrak{p}B\text{ is prime}.$
+[…]
 This proposition implies that if $A$ is reduced, and $\phi$ has prime extension property, then $B$ is also reduced.
 Indeed, if $A$ contains the zero ideal as a prime ideal, so does $B$.
-It also gives us an enormous class of examples.
-If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain, which is geometrically interesting.
-Before we give an example, we recall the following lemma.
+If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain.
+We are ready to give a geometric example if we recall one more lemma.
 Lemma 2.13.
-[Sta24, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
+[stacks-project, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
 Example 2.14.
@@
 The injectivity comes from the fact that both $y-x^{2}$ and $y-x^{2}-tx$ are irreducible in ${\mathbbm{k}}[t,x,y]$.
-We identify closed points of $A$ with elements of ${\mathbbm{k}}$, and obviously the $(t-c)$ in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension.
+We identify closed points of $A$ with elements of $\mathbbm{k}$}, and obviously the (tc)(t-c) in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension.
 The fiber $\operatorname{Spec}B_{(t-c)}$ is given by the plane curve $y=x^{2}+cx$ in $\mathbb{A}_{\mathbbm{k}}^{2}$, which is an irreducible variety, and those plane curves are even homeomorphic to each other.
-The situation doesn’t always perform so nicely.
-If we consider $A$ as ${\mathbbm{k}}[t,s]$ and $B$ as ${\mathbbm{k}}[t,s,x,y]/(y-x^{2}-tx)$, we also have $(t-c)$ as our point, then $B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ will be ${\mathbbm{k}}[s,s^{-1},x,y]/(y-x^{2}-cx)={\mathbbm{k}}(s)[x,y]/(y-x^{2}-cx)$.
-However we have $(t-a,s-b)$ as our point where $a$,$b\in{\mathbbm{k}}$, now the corresponding fiber is $\operatorname{Spec}{\mathbbm{k}}[x,y]/(y-x^{2}-ax)$, which is not homeomorphic to the previous one.
+[…]
 3. Resurgence of sum of ideals
-Reviewing the recent research of resurgence number of sums of ideals for polynomial rings over fields, our goal is to prove these results for sums of finitely generated ${\mathbbm{k}}$-algebras which are integral domains where ${\mathbbm{k}}$ is algebraically closed.
-In [BHJT21, Theorem 2.6] it has been proved that if $I$ and $J$ are ideals in different polynomial rings then
+First note that in this section, we assume that all rings and algebras are regular.
+We summarize current results of resurgence number of sums of ideals for polynomial rings over a field, our goal is to prove these results for sums of finitely generated graded $\mathbbm{k}$}-algebras which are integral domains, where $\mathbbm{k}$} is algebraically closed.
+In [fiberprojective, Theorem 2.6], it has been proved that if $I$ and $J$ are ideals in different polynomial rings over a same field, then
 $\rho_{a}(I+J)=\overline{\rho}_{\alpha}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}.$
-In [KNT24, Theorem 2.3], it has been proved that
+In [sharpbound, Theorem 2.3], it has been proved that
 $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
-In this section we prove similar results for finitely generated regular ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field.
-One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras.
-The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up.
-To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
-In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals.
-In this section, we assume that all things are regular.
+In our paper, [fiberprojective, Theorem 2.6] is generalized to Theorem 3.5, and [sharpbound, Theorem 2.3] is generalized to Theorem 3.6.
+The generalization is not trivial.
+First, it is obvious that the maps from two polynomials rings $A$ and $B$ over a field $\mathbbm{k}$} to $A\otimes_{{\mathbbm{k}}}B$ preserve resurgence numbers.
+However, for our generalization, we need to apply Corollary 2.10 to achieve that resurgence number of an ideal is same as resurgence number of its extension.
+Second, in the proof of Lemma 3.3, we use vector space method, while Gröbner basis method was used in the polynomial case but could not be used in $\mathbbm{k}$}-algebra case.
+In Lemma 3.1, we will specify that under proper conditions resurgence of the extension is the same as the resurgence of the ideals.
 Lemma 3.1.
 Let $A$ and $B$ be finitely generated $\mathbbm{k}$} algebras, where $\mathbbm{k}$} is algebraically closed, which are integral domains.
-If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes B$, then $\phi$ is flat and has prime extension property.
-Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$.
+If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes_{{\mathbbm{k}}}B$, then $\phi$ is flat and has prime extension property.
+Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$ for any proper ideal $I$ from $A$ (or $B$).
 Proof.
@@
 Observe that $R/\mathfrak{p}R=\left(A/\mathfrak{p}\right)\otimes_{\mathbbm{k}}B$ because $R$ is an integral domain.
-Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54].
+Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [algvar, Lemma 1.54].
 It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property.
-The last equality follows from Proposition 3.3.
+The last equality follows from Corollary 2.10.
 ∎
-Remark 4.2.
-To see why we need the conditions on $A$ and $B$, we will need some examples and theorems.
-Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring.
-For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring.
-Therefore, the condition being finitely generated is necessary.
-We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain.
-It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain.
-Also note by [TY03, Theorem 6(e)] $A\otimes B$ is regular.
-The next lemma also extends [BHJT21, Lemma 2.4].
-The proof of [BHJT21, Lemma 2.4] depends on divisibility in the polynomial ring.
+The next lemma also extends [fiberprojective, Lemma 2.4].
+The proof of [fiberprojective, Lemma 2.4] depends on divisibility in the polynomial ring.
 While working with general domains, we treated them as vector spaces and used properties of basis.
+Let us recall the following definition from [guardo2013asymptotic].
+Definition 3.2.
+Let $I$ be an ideal in a regular ring. Then $\lim_{t\to\infty}\sup\rho(I,t)$ is defined as follows:
+$\rho(I,t)=\sup\left\{\dfrac{m}{r}~|~I^{(m)}\not\subseteq I^{r},m\geqslant t,r\geqslant t\right\},\text{ and }\rho^{\lim\sup}_{a}(I)=\lim_{t\to\infty}\sup\rho(I,t).$
 Lemma 3.3.
@@
 Thus for any $h,r\in{\mathbb{N}}$ so that $h,r\geqslant t\geqslant s_{0}$ and $\dfrac{h}{r}\geqslant\theta$, one has $(I+J)^{r}\supset(I+J)^{(h)}\supset I^{(h)}$.
-Up to this point the proof idea is same as [BHJT21, Lemma 2.4].
+Up to this point the proof idea is same as [fiberprojective, Lemma 2.4].
 Now we need to show that $I^{(h)}\subseteq I^{r}$.
@@
 Let $W$ be a $\mathbbm{k}$}-basis of $J$, which extends to $W^{*}$, a $\mathbbm{k}$}-basis of $B$ containing $1$.
-Then $V\otimes W^{*}$ is a ${\mathbbm{k}}$-basis for $I\otimes B$ and $V^{*}\otimes W$ is a ${\mathbbm{k}}$-basis for $V\otimes W^{*}$.
-Also, $V^{*}\otimes W^{*}$ is a ${\mathbbm{k}}$-basis of $R=A\otimes B$.
+Then $V\otimes_{{\mathbbm{k}}}W^{*}$ is a $\mathbbm{k}$}-basis for $I\otimes_{{\mathbbm{k}}}B$ and $V^{*}\otimes_{{\mathbbm{k}}}W$ is a $\mathbbm{k}$}-basis for $V\otimes_{{\mathbbm{k}}}W^{*}$.
+Also, $V^{*}\otimes_{{\mathbbm{k}}}W^{*}$ is a $\mathbbm{k}$}-basis of $R=A\otimes_{{\mathbbm{k}}}B$.
 Let $f\in I^{(h)}$ then $f=\sum_{i=1}^{n}c_{i}v_{i}$, where $c_{i}\in{\mathbbm{k}}\setminus\{0\}$ and $v_{i}\in V$.
@@
 $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(I+J)^{(h)}\subset(I+J)^{r}=\sum_{j=0}^{r}(IR)^{i}(JR)^{r-j}=\sum_{j=0}^{r}\left((IR)^{j}\cap(JR)^{r-j}\right).$
-The last equality follows from [HNTT20, Lemma 3.1].
-Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes_{k}1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$.
+The last equality follows from [ha2020symbolic, Lemma 3.1].
+Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes 1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$.
 Since $1\notin W$ then $v\otimes 1\notin A\otimes J$.
-Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $11\leqslant j\leqslant r-$ 1.
+Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $1\leqslant j\leqslant r-1$.
 Hence $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(IR)^{h}$.
@@
 ∎
-The following generalizes [BHJT21, Lemma 2.5] to finitely generated algebra domains.
-The proof of [BHJT21, Lemma 2.5] depends on the binomial summation formula and numerical inequalities, which holds in general domains as well, so we omit the proof.
+The following generalizes [fiberprojective, Lemma 2.5] to finitely generated algebra domains.
+The proof of [fiberprojective, Lemma 2.5] depends on the binomial summation formula ([bionomial, Theorem 3.4]) and numerical inequalities, which holds in Noetherian algebras over a field as well, so we omit the proof.
 Lemma 3.4.
@@
 Proof.
-The proof has the same spirit as [BHJT21, Theorem 2.6] and it follows from Lemma 4.1, Lemma 4.3 and Lemma 4.4. ∎
-Now we give the result regarding the summation formula for resurgence of ideals in finitely generated ideals in ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed field.
+The proof has the same spirit as [fiberprojective, Theorem 2.6] and it follows from Lemma 3.1, Lemma 3.3 and Lemma 3.4. ∎
 Theorem 3.6.
 Let $A$ and $B$ be finitely generated algebras over an algebraically closed field $\mathbbm{k}$} which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals.
-Set $I+J=IR+JR$, where $R=A\otimes B$.
+Set $I+J=IR+JR$, where $R=A\otimes_{{\mathbbm{k}}}B$.
 Then
@@
 Proof.
-We give a sketch of the proof as it follows the same pattern as [KNT24, Theorem 2.3].
+We give a sketch of the proof as it follows the same pattern as [sharpbound, Theorem 2.3].
 First note that by Lemma 3.1, $\rho(IR)=\rho(I)$ and $\rho(JR)=\rho(J)$.
-The binomial summation formula for symbolic powers hold for ideals in noetherian ${\mathbbm{k}}$-algebras (see [HJKN23, Theorem 4.1].
-The other numerical inequalities in [KNT24, Theorem 2.3] follows straightforward.
+The binomial summation formula for symbolic powers hold for ideals in noetherian $\mathbbm{k}$}-algebras (see [ha2023binomial, Theorem 4.1].
+The other numerical inequalities in [sharpbound, Theorem 2.3] follows straightforward.
 Hence the required result follows.
 ∎
+Remark 3.7.
+To see why we need these conditions on $A$ and $B$ from Theorem 3.5, first note that, if $A$ (as a field extension over $\mathbbm{k}$}) is not finitely generated as a $\mathbbm{k}$}-algebra, then $A\otimes_{\mathbbm{k}}A$ may not a Noetherian ring by [vmos, Theorem 11].
+For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring.
+Therefore, the condition being finitely generated is necessary.
+We also need $A$ and $B$ to be integral domains as $\mathbbm{k}$}-algebras where $\mathbbm{k}$} is algebraically closed because of [algvar, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain.
+A famous example for violating the algebraically closed condition, then we do not have integral domain, is $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$.
+Moreover, note by [Regulartensor, Theorem 6(e)], $A\otimes_{{\mathbbm{k}}}B$ is regular if $A$ and $B$ are regular.